# Remove Symmetries to Control Model Expressivity

### v1

@@ Abstract
When symmetry is present in the loss function, the model is likely to be trapped in a low-capacity state that is sometimes known as a “collapse.” Being trapped in these low-capacity states can be a major obstacle to training across many scenarios where deep learning technology is applied. We first prove two concrete mechanisms through which symmetries lead to reduced capacities and ignored features during training. We then propose a simple and theoretically justified algorithm, syre, to remove almost all symmetry-induced low-capacity states in neural networks. The proposed method is shown to improve the training of neural networks in scenarios when this type of entrapment is especially a concern. A remarkable merit of the proposed method is that it is model-agnostic and does not require any knowledge of the symmetry.
@@ 1 Introduction
The unprecedented scale and complexity of modern neural networks, which incorporate a vast number of neurons and connections, inherently introduce a high degree of redundancy in model parameters. This complexity in the architecture and the design of loss functions often implies that the loss functions are invariant to various hidden, nonlinear, and nonlocal transformations of the model parameters. These invariant transformations, or “symmetries,” in the loss function have been extensively documented in the literature, with common examples including permutation, rescaling, scale, and rotation symmetries [1, 2, 3, 4, 5, 6, 7, 8, 9, 10].
A unifying perspective to understand how these symmetries affect learning is the framework of reflection symmetries [11]. [^1] Essentially, this is because the common symmetries in deep learning all contain $\mathbb{Z}_{2}$ as a subgroup. A per-sample loss function $\ell$ possesses a $P$-reflection symmetry if for all $\theta$, and all data points $x$,
$\ell((I-2P)\theta,x)=\ell(\theta,x).$ (1)
The solutions where $\theta=(I-P)\theta$ are the symmetric solutions with respect to $P$. It has been shown that almost all contemporary models contain quite a few reflection symmetries. Permutation, rescaling, scale, and rotation symmetries all imply the existence of one or multiple reflection symmetries in the loss.
Recent literature has shown that symmetries in the loss function of neural networks often lead to the formation of low-capacity saddle points within the loss landscape [12, 13, 14]. These saddle points are located at the symmetric solutions and often possess a lower capacity than the minimizers of the loss [11]. When a model encounters these saddle points during training, the model parameters are not only slow to escape them but also attracted to these solutions because these the gradient noise also vanish close to these saddles [14, 15]. Essentially, the model’s learning process stagnates, and it fails to achieve optimal performance due to reduced capacity. However, while many works have characterized the dynamical properties of training algorithms close to symmetric solutions, no methods are known to enable full escape from them.
@@
We introduce the notations and problem setting for our work in the next section. Closely related works are discussed in Section 3. All the proofs and experimental details are deferred to the appendix.
@@ 2 Discrete Symmetries in Neural Networks
For a matrix $A$, we use $A^{+}$ to denote the pseudo-inverse of $A$. For groups $U$ and $G$, $U\lhd G$ denotes that $U$ is a subgroup of $G$. For a vector $w$ and matrix $D$, $\|\theta\|^{2}_{D}:=w^{T}Dw$ is the norm of $w$ with respect to $D$. $\odot$ denotes the element-wise product between vectors.
@@
The second condition is due to the fact that there is a redundancy in the choice of $\theta^{\prime}$ when $\theta^{\prime}\neq 0$. Requiring $P\theta^{\prime}=\theta^{\prime}$ removes this redundancy and makes the choice of $\theta^{\prime}$ unique. Since every projection matrix can be written as a product of a (full-rank or low-rank) matrix $O$ with orthonormal columns, one can write $P=OO^{T}$ and refer to this symmetry as an $O$ symmetry. In common deep learning scenarios, it is almost always the case that $\theta^{\prime}=0$ (for example, this holds for the common cases of rescaling symmetries, (double) rotation symmetries, and permutation symmetries, see Theorem 2-4 of Ref. [11]). A consequence of $\theta^{\prime}=0$ is that the symmetric projection $P\theta$ of any $\theta$ always has a smaller norm than $\theta$: thus, a symmetric solution is coupled to the solutions of weight decay, which also favors small-norm solutions. As an example of reflection symmetry, consider a simple tanh network $f(\theta,x)=\theta_{1}\tanh(\theta_{2}x)$. The model output is invariant to a simultaneous sign flip of $\theta_{1}$ and $\theta_{2}$. This corresponds to a reflection symmetry whose projection matrix is the identity $P=((1,0),(0,1))$. The symmetric solutions correspond to the trivial state where $\theta_{1}=\theta_{2}=0$.
@@
Therefore, requiring $P\theta^{\prime}=\theta^{\prime}$ reduces different manifestations of the symmetry to a unique one and simply shifting the function will not remove any symmetry. This proposition emphasizes the subtle difficulty in removing a symmetry.
@@ 3 Related Works
One closely related work is Ref. [11], which shows that every reflection symmetry in the model leads to a low-capacity solution that is energetically favored when weight decay is used. This is because the minimizer of the weight decay is coupled with stationary points of the reflection symmetries – the projection of any parameter to a symmetric subspace always decreases the norm of the parameter, and is thus energetically preferred by weight decay. Our work develops a method to decouple symmetries and weight decay, thus avoiding collapsing into low-capacity states during training. Besides weight decay, an alternative mechanism for this type of capacity loss is gradient-noise induced collapse, which happens when the learning rate - batchsize ratio is high [14, 15].
Contemporarily, Ref. [16] empirically explores how removing symmetries can benefit neural network training and suggests a heuristic for removing symmetries by hold a fraction of the weights unchanged during training. However, the proposed method can only remove permutation symmetries in fully connected layers. This is particularly a problem because most of the symmetries in nonlinear systems are unknown and hidden [17, 18]. In sharp contrast, the technique proposed in our work is both architecture-independent and symmetry-agnostic and provably removes all known and unknown $P$-reflection symmetries in the loss.
@@ 4 Symmetry Impairs Model Capacity
We first show that reflection symmetry directly affects the model capacity. For simplicity, we let $\theta^{\prime}=0$ for all symmetries. Let $f(x,\theta)\in\mathbb{R}$ be a Taylor-expandable model that contains a $P$-reflection symmetry. Let $\Delta=\theta-\theta_{0}$. Then, close to any symmetric point $\theta_{0}$ (any $\theta_{0}$ for which $P\theta_{0}=0$), for all $x$, Ref. [11] showed that
$f(x,\theta)-f(x,\theta_{0})=$ $\underbrace{\nabla_{\theta}f(x,\theta_{0})(I-P)\Delta+O(\|P\Delta\|)^{2}}_{\text{Symmetry subspace}}+\underbrace{\frac{1}{2}\Delta^{T}PH(x)P\Delta+O(\|\Delta\|^{3})}_{\text{Symmetry-broken subspace}},$ (2)
@@
$\theta^{*}=A^{+}\sum_{x}\nabla_{\theta}f(x,\theta_{0})^{T}y(x)$ (3)
@@
This means that in the kernel regime [^2] Technically, this is the lazy training limit [19]., being at a symmetric solution implies that the feature kernel features are being masked by the projection matrix:
@@
The existence of this type of dimension reduction when symmetry is present has also been noticed by previous works in case of permutation symmetry [1]. Intuitively, the above two results follow from the fact that the symmetric subspaces of reflection symmetries (and any general discrete symmetries) are a linear subspace, and so gradient descent (and thus gradient flow) cannot take the model away from it, despite the discretization error. This means that, essentially, the model is identical to a model with a strictly smaller dimension throughout its training, no matter whether the training is through GD or SGD. When there are multiple symmetries, there is a compounding effect. An alternative perspective to this problem is through the classical framework of singular learning theories, where each symmetry corresponds to a singularity and directly affects the asymptotic performance of the model [20].
See Figure 1 for an illustration. We initialize a two-layer ReLU neural network on a low-capacity state where a fraction of the hidden neurons are identical (corresponding to the symmetric states of the permutation symmetry [11]) and train with and without removing the symmetries. We see that when the symmetries are removed (with the method proposed in the next section), the model is no longer stuck at these neuron-collapsed solutions.
@@ 5.1 Countable Symmetries
We seek an algorithm that eliminates the reflection symmetries from the loss function $\ell$. We show that when the number of reflection symmetries in the loss function is finite, one can completely remove them using a simple technique. The symmetries are required to have the following property and the loss function is assumed to obey assumption 1.
@@
Let $\ell$ satisfy Property 1 and Assumption 1. Then, with probability $1$ (over the sampling of $\theta_{0}$), there exists no projection matrix $P$ and reflection point $\theta^{\prime}$ such that $\ell_{\rm r}$ has the $(\theta^{\prime},P)$-symmetry.
The core mechanism of this theorem is decoupling of the solutions of the symmetries from the solutions of the weight decay. With weight decay, a solution with a small norm is favored, whereas with a random bias, the symmetric solutions are shifted by a small constant and no longer overlap with solutions that have a small norm. Equivalently, this method can be implemented as a random bias in the $L_{2}$ regularization: $\ell_{\rm r}(\theta,x)=\ell(\theta)+\gamma\|\theta+\theta_{0}\|^{2}$. This is a result of the fact that simple translation does not change the nature of the function. There are two remarkable parts of the theorem: (1) it not only removes all existing symmetries but also guarantees that there are no remaining ones; (2) the proof works as long as $\theta_{0}$ is nonzero. This hints at the possibility of using a small and random $\theta_{0}$, which removes all symmetries in principle and also does not essentially affect the solutions of the original objective. In this work, we will refer to the method in Eq. (6) as syre, an abbreviation for “symmetry removal.”
@@ 5.2 Uncountably Many Symmetries
In deep learning, it is possible for the model to simultaneously contain infinitely many reflection symmetries. This happens, for example, when the model parameters have the rotation symmetry or the double rotation symmetry (common in self-supervised learning problems or transformers). It turns out that simply adding a static bias and weight decay is not sufficient to remove all symmetries in this case.
@@ 5.3 Strength of Symmetry Removal
$\forall g,\ \ell(\theta)=\ell(g\theta).$ (10)
@@
The following theorem shows that at every symmetric solution, there exists an escape direction with a strong gradient that pulls the parameters from every subgroup of the related symmetry. In other words, it is no longer possible to be trapped in a symmetric solution. While general groups appear less common in deep learning scenarios, they exist in a lot of application scenarios with equivariant networks [21], where it is common to incorporate generic group structures into the model.
@@
Note that $\overline{U}$ is a projection matrix: $\overline{U}\,\overline{U}=\overline{U}$. This means that $I-\overline{U}$ is also a projection matrix. Importantly, $\overline{U}$ projects a vector into the symmetric subspace. For any $u\in U$, $u\overline{U}=\overline{U}$. Likewise, $I-\overline{U}$ projects any vector into the symmetry-broken subspace, a well-known result in the theory of finite groups [22]. We denote by
@@
    if $G$ is abelian, $\min_{V\lhd U}\Delta_{V}=\Omega(\gamma\sigma_{0}{\rm rank}(I-\overline{V})|\Gamma(U)|^{-1})$; [^5] This result can be generalized to the case where all projectors $\bar{V}$ commute with each other, even if $G$ is nonabelian. Essentially, this result is a consequence of the properties of the representations of $G$ and not due to $G$ itself.
@@
Item (2) of the theorem is essentially due to the fact that removing symmetries from a larger group can be reduced to removing them from one of its subgroups. Conceptually, this result has the same root as the classical theory of combinatorial designs, where averaging over a subset of groups has the same effect as averaging over the whole group [23]. In general, $1\leq Z\leq|G|$ (and sometimes $\ll|G|$), and so this scaling is not bad. For the part (3) of the theorem, the term $|\Gamma(U)|$ is especially meaningful. It is well-known that $|\Gamma(U)|\leq\log|U|$, and so the worst-case symmetry-breaking strength is only of order $1/\log|U|$, which is far slower than what one would expect. In fact, for a finite group with size $N$, the number of subgroups can grow as fast as $N^{\log N}$ [24], and thus, one might naively think that the minimal breaking strength decreases as $N^{-\log N}$. This theorem shows that the proposed method is highly effective at breaking the symmetries in the loss function or the model.
Figure 2: With the proposed method, the robustness of the feature increases with the input dimension. The same may happen under vanilla training, but the effect is not strong. The covariance matrix is measured from the first-layer output with batch size $1000$, and we set eigenvalues larger than $10^{-4}$ to zero.
@@
Role of Noise.
A core concept in our theory is that random noise, even if static, can help remove the symmetries. A related question is whether simply having noise in the input or label is sufficient, and so no static bias is required. The answer is no. In Appendix B.1, we find that adding input noise does help alleviate collapse, whereas adding label noise exacerbates it.

### v3

@@ Abstract
When symmetry is present in the loss function, the model is likely to be trapped in a low-capacity state that is sometimes known as a “collapse.” Being trapped in these low-capacity states can be a major obstacle to training across many scenarios where deep learning technology is applied. We first prove two concrete mechanisms through which symmetries lead to reduced capacities and ignored features during training and inference. We then propose a simple and theoretically justified algorithm, syre, to remove almost all symmetry-induced low-capacity states in neural networks. When this type of entrapment is especially a concern, removing symmetries with the proposed method is shown to correlate well with improved optimization or performance. A remarkable merit of the proposed method is that it is model-agnostic and does not require any knowledge of the symmetry.
@@ 1 Introduction
The unprecedented scale and complexity of modern neural networks, which contain a vast number of neurons and connections, inherently introduce a high degree of redundancy in model parameters. This complexity in the architecture and the design of loss functions often implies that the loss functions are invariant to various hidden and nonlinear transformations of the model parameters. These invariant transformations, or “symmetries,” in the loss function have been extensively documented in the literature, with common examples including permutation, rescaling, scale, and rotation symmetries (Simsek et al., 2021; Entezari et al., 2021; Dinh et al., 2017; Neyshabur et al., 2014; Tibshirani, 2021; Zhao et al., 2023; Godfrey et al., 2022; Ziyin et al., 2024).
A unifying perspective to understand how these symmetries affect learning is the framework of reflection symmetries (Ziyin, 2024). [^1] Essentially, this is because the common symmetries in deep learning all contain $\mathbb{Z}_{2}$ as a subgroup. A per-sample loss function $\ell$ possesses a $P$-reflection symmetry if for all $\theta$, and all data points $x$,
$\vspace{-1mm}\ell((I-2P)\theta,x)=\ell(\theta,x).$ (1)
The solutions where $\theta=(I-P)\theta$ are the symmetric solutions with respect to $P$. It has been shown that almost all models contain quite a few reflection symmetries. Permutation, rescaling, scale, and rotation symmetries all imply the existence of one or multiple reflection symmetries in the loss.
Recent literature has shown that symmetries in the loss function of neural networks often lead to the formation of low-capacity saddle points within the loss landscape (Fukumizu, 1996; Li et al., 2019). These saddle points are located at the symmetric solutions and often possess a lower capacity than the minimizers of the loss. When a model encounters these saddle points during training, the model parameters are not only slow to escape them but also attracted to these solutions because these the gradient noise also vanish close to these saddles (Chen et al., 2023). Essentially, the model’s learning process stagnates, and it fails to achieve optimal performance due to reduced capacity. However, while many works have characterized the dynamical properties of training algorithms close to symmetric solutions, no methods are known to enable full escape from them.
@@
We introduce the notations and problem setting for our work in the next section. Closely related works are discussed in Section 3. All proofs and experimental details are deferred to the appendix.
@@ 2 Discrete Symmetries in Neural Networks
For a matrix $A$, we use $A^{+}$ to denote the pseudo-inverse of $A$. For groups $U$ and $G$, $U\lhd G$ denotes that $U$ is a subgroup of $G$. For a vector $\theta$ and matrix $D$, $\|\theta\|^{2}_{D}:=\theta^{T}D\theta$ is the norm of $\theta$ with respect to $D$. $\odot$ denotes the element-wise product between vectors.
@@
The second condition is due to the fact that there is a redundancy in the choice of $\theta^{\prime}$ when $\theta^{\prime}\neq 0$. Requiring $P\theta^{\prime}=\theta^{\prime}$ removes this redundancy and makes the choice of $\theta^{\prime}$ unique. Since every projection matrix can be written as a product of a (full-rank or low-rank) matrix $O$ with orthonormal columns, one can write $P=OO^{T}$ and refer to this symmetry as an $O$ symmetry. In common deep learning scenarios, it is almost always the case that $\theta^{\prime}=0$ (for example, this holds for the common cases of rescaling symmetries, (double) rotation symmetries, and permutation symmetries, see Theorem 2-4 of Ziyin (2024)). A consequence of $\theta^{\prime}=0$ is that the symmetric projection $P\theta$ of any $\theta$ always has a smaller norm than $\theta$: thus, a symmetric solution is coupled to the solutions of weight decay, which also favors small-norm solutions. As an example of reflection symmetry, consider a simple tanh network $f(\theta,x)=\theta_{1}\tanh(\theta_{2}x)$. The model output is invariant to a simultaneous sign flip of $\theta_{1}$ and $\theta_{2}$. This corresponds to a reflection symmetry whose projection matrix is the identity $P=((1,0),(0,1))$. The symmetric solutions correspond to the trivial state where $\theta_{1}=\theta_{2}=0$.
@@
Therefore, requiring $P\theta^{\prime}=\theta^{\prime}$ is without loss of generality: it only reduces different manifestations of the symmetry to a unique one and simply shifting the function will not remove any symmetry. This proposition emphasizes the subtle difficulty in removing a symmetry.
@@ 3 Related Works
One closely related work is that of Ziyin (2024), which shows that every reflection symmetry in the model leads to a low-capacity solution that is favored when weight decay is used. This is because the minimizer of the weight decay is coupled with stationary points of the reflection symmetries – the projection of any parameter to a symmetric subspace always decreases the norm of the parameter, and is thus energetically preferred by weight decay. Our work develops a method to decouple symmetries and weight decay, thus avoiding collapsing into low-capacity states during training. Besides weight decay, an alternative mechanism for this type of capacity loss is gradient-noise induced collapse, which happens when the learning rate - batchsize ratio is high (Chen et al., 2023).
Contemporarily, Lim et al. (2024) empirically explores how removing symmetries can benefit neural network training and suggests a heuristic for removing symmetries by hold a fraction of the weights unchanged during training. However, the proposed method is only proved to work for explicit permutation symmetries in fully connected layers. This is particularly a problem because most of the symmetries in nonlinear systems are unknown and hidden (Cariglia, 2014; Frolov et al., 2017). In sharp contrast, the technique proposed in our work is both architecture-independent and symmetry-agnostic and provably removes all known and unknown discrete symmetries in the loss.
@@ 4 Symmetry Impairs Model Capacity
We first show that reflection symmetry directly affects the model capacity. For simplicity, we let $\theta^{\prime}=0$ for all symmetries. Let $f(x,\theta)\in\mathbb{R}$ be a Taylor-expandable model that contains a $P$-reflection symmetry. Let $\Delta=\theta-\theta_{0}$. Then, close to any symmetric point $\theta_{0}$ (any $\theta_{0}$ for which $P\theta_{0}=0$), for all $x$, Ziyin (2024) showed that
$f(x,\theta)-f(x,\theta_{0})=$ $\underbrace{\nabla_{\theta}f(x,\theta_{0})(I-P)\Delta+O(\|P\Delta\|)^{2}}_{\text{Symmetry subspace}}+\underbrace{\frac{1}{2}\Delta^{T}PH(x)P\Delta+O(\|\Delta\|^{3})}_{\text{Symmetry-broken subspace}},\vspace{-2mm}$ (2)
@@
$\theta^{*}=A^{+}\sum_{x}\nabla_{\theta}f(x,\theta_{0})^{T}y(x)\vspace{-1mm}$ (3)
@@
This means that in the kernel regime [^2] Technically, this is the lazy training limit (Chizat et al., 2018)., being at a symmetric solution implies that the feature kernel features are being masked by the projection matrix:
@@
The existence of this type of dimension reduction when symmetry is present has also been noticed by previous works in case of permutation symmetry (Simsek et al., 2021). Intuitively, the above two results follow from the fact that the symmetric subspaces of reflection symmetries (and any general discrete symmetries) are a linear subspace, and so gradient descent (and thus gradient flow) cannot take the model away from it, despite the discretization error. This means that, essentially, the model is identical to a model with a strictly smaller dimension throughout its training, no matter whether the training is through GD or SGD. When there are multiple symmetries, there is a compounding effect. See Figure 1 for an illustration. We initialize a two-layer ReLU neural network on a low-capacity state where a fraction of the hidden neurons are identical (corresponding to the symmetric states of the permutation symmetry) and train with and without removing the symmetries. We see that when the symmetries are removed (with the method proposed in the next section), the model is no longer stuck at these neuron-collapsed solutions. Also, see Section C for the evolution of NTK and gradient covariance during training.
@@ 5.1 Countable Symmetries
We seek an algorithm that eliminates the reflection symmetries from the loss function $\ell$. This subsection will show that when the number of reflection symmetries in the loss function is countable, one can completely remove them using a simple technique. The symmetries are required to have the following property and the loss function is assumed to obey assumption 1.
@@
Let $\ell$ satisfy Property 1 and Assumption 1. Then, with probability $1$ over the sampling of $\theta_{0}$, there exists no projection matrix $P$ and reflection point $\theta^{\prime}$ such that $\ell_{\rm r}$ has the $(\theta^{\prime},P)$-symmetry.
The core mechanism of this theorem is decoupling of the solutions of the symmetries from the solutions of the weight decay. With weight decay, a solution with a small norm is favored, whereas with a random bias, the symmetric solutions are shifted by a small constant and no longer overlap with solutions that have a small norm. Equivalently, this method can be implemented as a random bias in the $L_{2}$ regularization: $\ell_{\rm r}(\theta,x)=\ell(\theta)+\gamma\|\theta+\theta_{0}\|^{2}$. This is a result of the fact that simple translation does not change the nature of the function. There are two remarkable parts of the theorem: (1) it not only removes all existing symmetries but also guarantees that there are no remaining ones; (2) the proof works as long as the variance of $\theta_{0}$ is nonzero. This hints at the possibility of using a small and random $\theta_{0}$, which removes all symmetries in principle and also does not essentially affect the solutions of the original objective. In this work, we will refer to the method in Eq. (6) as syre, an abbreviation for “symmetry removal.”
@@ 5.2 Uncountably Many Symmetries
In deep learning, it is possible for the model to simultaneously contain infinitely many reflection symmetries. This happens, for example, when the model parameters have the rotation symmetry or the double rotation symmetry (common in self-supervised learning problems Ziyin et al. (2023) or transformers). It turns out that adding a static bias and weight decay is not sufficient to remove all symmetries, but we will show that a simple modification would suffice.
@@ 5.3 Strength of Symmetry Removal
$\forall g\in G,\ \ell(\theta)=\ell(g\theta).$ (10)
@@
The following theorem shows that at every symmetric solution, there exists an escape direction with a strong gradient that pulls the parameters away from every subgroup of the related symmetry. In other words, it is no longer possible to be trapped in a symmetric solution. While general groups appear less common, they exist in a lot of application scenarios with equivariant networks (Cohen & Welling, 2016), where it is common to incorporate group structures into the model.
@@
Note that $\overline{U}$ is a projection matrix: $\overline{U}\,\overline{U}=\overline{U}$. This means that $I-\overline{U}$ is also a projection matrix. Importantly, $\overline{U}$ projects a vector into the symmetric subspace. For any $u\in U$, $u\overline{U}=\overline{U}$. Likewise, $I-\overline{U}$ projects any vector into the symmetry-broken subspace, a well-known result in the theory of finite groups (Gorenstein, 2007). We denote by
@@
    if $G$ is abelian, $\min_{V\lhd U}\Delta_{V}=\Omega(\gamma\sigma_{0}{\rm rank}(I-\overline{V})|\Gamma(U)|^{-1})$; [^5] This result can be generalized to the case where all projectors $\bar{V}$ commute with each other, even if $G$ is nonabelian. Namely, this is a consequence of the properties of the representations of $G$ and of $G$ itself.
@@
Figure 2: With the proposed method, the rank of the feature covariance matrix increases with the input dimension. The same may happen under vanilla weight decay, but the effect is not strong. The covariance is computed from the first-layer output with batch size $1000$, and eigenvalues larger than $10^{-4}$ are truncated to zero.
Item (2) of the theorem is essentially due to the fact that removing symmetries from a larger group can be reduced to removing them from one of its subgroups. Conceptually, this result has the same root as the classical theory of combinatorial designs, where averaging over a subset of groups has the same effect as averaging over the whole group (Lindner & Rodger, 2017). In general, $1\leq Z\leq|G|$ (and sometimes $\ll|G|$), and so this scaling is not bad. For the part (3) of the theorem, the term $|\Gamma(U)|$ is especially meaningful. It is well-known that $|\Gamma(U)|\leq\log|U|$, and so the worst-case symmetry-breaking strength is only of order $1/\log|U|$, which is far slower than what one would expect. In fact, for a finite group with size $N$, the number of subgroups can grow as fast as $N^{\log N}$ (Borovik et al., 1996), and thus, one might naively think that the minimal breaking strength decreases as $N^{-\log N}$. This theorem shows that the proposed method is highly effective at breaking the symmetries in the loss function or the model.